OpenStudy (anonymous):

HELPPPPPPPPP!!!!! If you hit your brakes 137.4 ft away from a camel, and decelerate at 28 ft/sec^2, will you stop before hitting the camel? Show work please!

OpenStudy (anonymous):

.5*a*t^2

OpenStudy (anonymous):

how fast is the car going

OpenStudy (dumbcow):

what was speed before hitting brakes

OpenStudy (anonymous):

Cal is driving a car at the speed of 60 mph (88 ft/sec), when suddenly he sees a camel standing on the road 199 ft. in front of him. Unable to swerve, he slams on his brakes. If he has a reaction time of .7 sec. before applying the brake & the car decelerates at a constant rate of 28 ft/sec^2, will he stop before hitting the camel?

OpenStudy (dumbcow):

distance equation \[d = \frac{1}{2}at^{2} + v_0 t \] a = -28 v_0 = 88 if d<137.4 then car stops in time Now to find t, use velocity equation \[v = at +v_0\] set it to 0, solve for t \[\rightarrow t = \frac{-v_0}{a}\] plug this into distance equation for t to see how far car travels in that time

OpenStudy (anonymous):

for the velocity equation, what to i plug in for velocity, 88? and initial velocity on the velocity formula, 137.4??

OpenStudy (dumbcow):

in both equations its 88 for initial velocity

OpenStudy (anonymous):

okay, but then for v=at+v 0, i plug 88 for v0 and what do i plug in for v because im solving for t?

OpenStudy (dumbcow):

oh sorry you are making v=0 ....when you stop your velocity is 0 thats how i came up with value for t

OpenStudy (anonymous):

oh okay! thanks so much! :)

OpenStudy (dumbcow):

your welcome...looks like he hits the camel :)